In developing patient appointment schedules, a medical center wants to estimate the mean time that a staff member spends with each patient. How large a sample should be taken if the desired margin of error is two minutes at a 95% level of confidence? How large a sample should be taken for a 99% level of confidence? Use a planning value for the population standard deviation of 9 minutes.

95% Confidence (to the nearest whole number):

99% Confidence (to the nearest whole number):
Answer #1


Margin of error is 2 minutes that is

Margin of error=Critical Value* SE of the statistic=2

Critical Value of t for 95% interval= 2 (for sample size>30)

2=2*SD/Sqrt(Sample Size)

Sqrt(Sample Size)=SD=9

Sample Size=81

For 99% Confidence

Critical Value of t for 99% interval=2.62

Margin of Error =Critical Value*SE of the statistic =2

2=2.62*9/Sqrt(Sample Size)

2/(2.62*9)=1/(sqrt(sample Size))

1/0.048=sqrt(sample size)

11.792=sqrt(Sample Size)

Sample Size=139.06=140 approximately

Hence the answers are as follows

95% CI for 2 minutes of Margin of error we need Sample size of at least 81

99% CI for 2 minutes of Margin of error we need Sample size of at least 140